Data from a weather forecast station indicate that the annual precipitation in Bandsville is a normal random variable with a mean of 40 inches and a standard deviation of 10 inches. Find the probability that the total precipitation during the next 2 years will exceed 86 inches.
Answer #1

Suppose, random variable \tiny X_i\left ( i=1,2 \right ) denotes amount (in inches) of rain in i-th year.

\tiny \therefore X_i\sim N\left(\mu=40,\sigma^2=10^2\right)

We define,

\tiny Y=X_1+X_2

\tiny \therefore E\left (Y \right )=E\left (X_1+X_2 \right )=E\left (X_1 \right )+E\left (X_2 \right )=40+40=80

\tiny \therefore Var\left (Y \right )=Var\left (X_1+X_2 \right )=Var\left (X_1 \right )+Var\left (X_2 \right )=10^2+10^2=14.14214^2

\tiny \therefore Y\sim N\left(\mu_y=80,\sigma_y^2=14.14214^2\right)

Required probability is given by

\tiny P\left ( Y>86 \right )=P\left ( \frac{Y-\mu_y}{\sigma_y}>\frac{86-80}{14.14214} \right )

\tiny =P\left ( Z>0.4242639\right )=0.3356867 [Using R-code '1-pnorm(0.4242639)']